OpenStudy (anonymous):

solve 3x=69 round answer to nearest hundredth

OpenStudy (anonymous):

x= 69/3

OpenStudy (anonymous):

options are x=4.89 x=2.13 x=3.85 x=4.23

OpenStudy (anonymous):

\[3x=69\]Divide both sides by 3.\[x=\frac{69}{3}=23\]

OpenStudy (anonymous):

I think you incorrectly gave the question then.

OpenStudy (anonymous):

3^x=69

OpenStudy (anonymous):

There you go. Haha

OpenStudy (anonymous):

So the inverse of that is going to be logarithmic.

OpenStudy (anonymous):

\[3^x=69\]Take the logarithm (base 3) of both sides.\[\log_3(3^x)=\log_3(69)\]log3 and 3 cancel out.\[x=\log_3(69)=3.85\]

OpenStudy (anonymous):

ohh! alright, thank yall!

OpenStudy (anonymous):

You're welcome